# MonoPatchNeRF: Improving Neural Radiance Fields with Patch-based Monocular Guidance

### v1

@@ Abstract
The latest regularized Neural Radiance Field (NeRF) approaches produce poor geometry and view extrapolation for multiview stereo (MVS) benchmarks such as ETH3D. In this paper, we aim to create 3D models that provide accurate geometry and view synthesis, partially closing the large geometric performance gap between NeRF and traditional MVS methods. We propose a patch-based approach that effectively leverages monocular surface normal and relative depth predictions. The patch-based ray sampling also enables the appearance regularization of normalized cross-correlation (NCC) and structural similarity (SSIM) between randomly sampled virtual and training views. We further show that “density restrictions” based on sparse structure-from-motion points can help greatly improve geometric accuracy with a slight drop in novel view synthesis metrics. Our experiments show 4x the performance of RegNeRF and 8x that of FreeNeRF on average F1@2cm for ETH3D MVS benchmark, suggesting a fruitful research direction to improve the geometric accuracy of NeRF-based models, and sheds light on a potential future approach to enable NeRF-based optimization to eventually outperform traditional MVS.
@@
Figure 1: We present MonoPatchNeRF on a sparse-view scene facade. Our method renders realistic images and accurate normals from the test view and reconstructs the complete mesh.
@@ 1 Introduction
Modeling 3D scenes from imagery is a core problem in computer vision, directly useful for mapping, facility assessment, robotics, construction monitoring, and more. These applications require both accurate geometry and realistic visualization from novel views, and may involve large scenes with sparse views.
Neural Radiance Field (NeRF) [23] approaches can synthesize novel views with high fidelity, especially when interpolating training views. The NeRF approach also has great potential for 3D geometry estimation, as a single model can be optimized according to many losses and constraints, encoding photometric consistency, single-view predictions, and material properties. However, this potential is currently unrealized — even the latest regularized approaches like RegNeRF [24] and FreeNeRF [34] produce poor geometry and view extrapolation for multiview stereo (MVS) benchmarks such as ETH3D [29]. While MVS methods typically produce superior geometry in numerous contexts, they often result in noisy and incomplete models, and they do not support high-quality rendering. Furthermore, enhancing MVS methods by introducing additional cues and priors is challenging due to the inherent heuristics and algorithmic complexity.
Recent neural field based works have led to some improvements in geometric fidelity and the capability to handle sparse views [24, 40, 10, 34]. MonoSDF [38] incorporates monocular surface normal and depth predictions in a signed distance function (SDF) model. This works well for objects or small indoor scenes, but due to the initialization with a sphere, smoothness priors, and volumetric representation, this approach does not readily provide good results for large scenes, open scenes, or those with fine structures like bicycle racks. In more recent SDF based work, Neuralangelo [16], smoothness priors are learned as a curvature loss with coarse-to-fine scheduled learning, which works well on challenging thin structures, but our experiments indicate limited effectiveness in a sparse view setup. RegNerf [24] incorporates losses of appearance likelihood and smooth depth for patches on randomly sampled virtual views, improving results when sparse input views are available. Yet, as our experiments show, this virtual view patch-based regularization is not sufficient to achieve geometrically accurate models. While not complete solutions, these approaches contain two critical ideas that we will build on: (1) leverage monocular surface normal and depth predictions to guide the model; (2) encourage virtual view appearance consistency to further regularize the model.
In this paper, we aim to create 3D models that provide accurate geometry and view synthesis, using a patch-based approach that effectively leverages monocular depth/normal predictions and virtual view appearance consistency priors. We have found that density-based models are better suited for encoding details in large scenes compared to SDF-based models. However, integrating monocular predictions of surface normal and up-to-scale depth into the ray-based sampling of density models is difficult because surface normal and relative depth are not well-defined along a ray in the density model. With our patch-based sampling technique, we use these monocular cues more effectively. Additionally, patch-based sampling allows us to take advantage of structural losses for regularization, such as normalized cross-correlation (NCC) and structural similarity (SSIM). These are robust to varying lighting conditions, hence extensively employed in classic MVS. Using these losses, we boost patch-level photometric consistency across randomly sampled virtual views, significantly constrain the 3D structure, and reduce artifacts. We further leverage monocular depth predictions by translating and scaling them based on sparse structure-from-motion (SfM) points and restricting surfaces to lie within a dilation of the union of all thus aligned points. This “density restriction” greatly improves geometric accuracy, with a slight drop in novel view synthesis metrics. With these improvements, we achieve promising results on the ETH3D benchmark, outperforming other NeRF-based methods in novel view synthesis and by a large margin in geometry. Results fall well short of state-of-art MVS methods for geometric accuracy, but we believe this work supports the potential of NeRF-based optimization to eventually surpass traditional MVS approaches.
In summary, this paper offers the following contributions:
  • We demonstrate the effective use of monocular surface normal and relative depth predictions for density-based NeRF models, proposing losses based on patch-based ray sampling optimization and occupancy constraints.
  • Using patch-based ray sampling, we demonstrate the effectiveness of NCC and SSIM photometric consistency losses between patches from virtual views and training views, creating stronger constraints than the appearance likelihood and depth smoothness constraints of RegNeRF.
  • Our method significantly improves geometric accuracy, achieving 4x the performance of RegNeRF and 8x that of FreeNeRF on average F1@2cm for the ETH3D MVS benchmark and produces competitive results in novel view synthesis, ranking best in terms of SSIM and LPIPS. These findings indicate a fruitful research direction to improve the geometric accuracy of NeRF-based models.
@@ 2 Related Works
Reconstructing the 3D scene from a set of images is a fundamental problem in computer vision. Multi-view Stereo (MVS), aiming to reconstruct geometry given images with known poses, is a well-studied field and develops from early works [18] with pure photometric scoring to more recent approaches that incorporate deep learning techniques [35, 36, 9, 13]. While there are various formulations of scene representation, state-of-the-art methods typically predict the depth map of each image based on photometric consistency with a set of source views [28, 12, 19, 13], and use geometric consistency across views to fuse the depths together into a single point cloud [28, 6]. The fused point clouds are evaluated against the ground truth geometry for precision-and-recall [11], or accuracy-and-completeness [29], with $F_{1}$ score combining the harmonic mean of the two values. Different from MVS, our method constructs a volumetric density representation and extracts depth maps using volumetric rendering of expected depth at source views, followed by a conventional depth-map fusion pipeline [6].
While MVS approaches solve for precise geometries [11, 29], the novel view rendering of meshes or point clouds generated by MVS is somewhat inferior compared to Neural Radiance Field (NeRF), a recently popular alternative for 3d reconstruction. NeRF [23] provides an elegant and effective solution for novel view synthesis. While NeRF has been shown to reliably estimate geometry and appearance in dense captures, multiple papers [37, 10, 26, 24, 34] show that it often fails to converge correctly in outward-facing, sparse, or wide-baseline inputs. PixelNeRF. [37] and DietNeRF [10] propose learning based feed-forward solutions that utilize prior knowledge for sparse input scenario. PixelNeRF represents the scene as extracted CNN features, and DietNeRF samples random poses and supervises random views using consistency with CLIP [26]. Both methods manage to render smooth novel views for scenes with sparse inputs, but suffer from degradation in rendering quality due to the low-resolution CNN features or indirect supervision.
Extending on PixelNeRF and DietNeRF, RegNeRF [24] proposes to use stronger regularization on randomly sampled patches via a geometric smoothness loss and an adversarial photometric loss. SPARF [30] jointly optimizes the NeRF models and camera poses with extracted pixels matches on input views, and improves performance given sparse inputs. Meanwhile, FreeNeRF [34] which modulates frequency in positional-encoded space shows that a simple coarse-to-fine approach is able to alleviate object-centric data-sparsity [1] and front-facing [21] scenes. DeLiRa [7] samples synthesized view points and warps the dilated rays for the photometric consistency loss. Previous methods target at solving the sparse views problem with more regularization through learned prior or virtual view smoothing geometry, but do not fully utilize the underlying geometry consistency. Our work extends RegNeRF [24] for sampling virtual views, but apply a photometric loss between virtual and training patches through learned geometry consistency for a better per scene regularization.
Using image-based priors in regularization has also been shown effective in training Neural SDF and NeRF based models. Roessle et al. [27] and NerfingMVS [32] use a pretrained network to estimate dense depth given sparse SfM points, and supervise the NeRF model with the estimated depth. However, the sparse depth from SfM often contains noise that can be passed down to the dense prediction [3, 25]. MonoSDF [38] proposes to use depth and normals estimated with RGB images and significantly improve the geometric quality of SDF models. NeuralWarp [4] proposes cross view photometric consistency for training patches as MVS with visibility information from SfM. While also using patch warping during optimization, our method enforces a consistency between virtual patches and training patches, which fits better for sparse view scenario, and we apply occlusion mask with geometry consistency determined by network outputs, as virtual patches do not have the visbility prior (Figure 3). Though with better geometry, SDF based approaches do not perform well in inferring high-frequency details such as bicycles and chair arms. Neuralangelo [16] partly resolves this problem using coarse-to-fine optimization and improves the geometry by using numerical gradient for higher derivatives, but our experiments indicate limited effectiveness in a sparse view scenario. While monocular cues have been widely applied, it is well known that the monocular depth is consistent locally, and the global inconsistency issue can lead to a worse geometry. We instead incorporate patch based monocular priors with a local scale-and-shift transformation, to better utilize the strength of monocular depth.
@@ 3 Method
Given a collection of posed images capturing a large unbounded scene, our goal is to construct a 3D neural radiance field representation that not only renders high-quality images from any viewpoint but also provides accurate and complete surface geometry. We achieve this by introducing three novel components: 1) patch-based monocular geometry cues, allowing more accurate and structured geometry; 2) patch-level photometric consistency regularization across virtual and training viewpoints, enhancing the overall rendering and geometry quality 3) density restriction through monocular cues and sparse geometry, inhibiting NeRF from predicting opacity on obvious empty space. Figure 2 provides a summary of our approach.
3.1 Background: Neural Radiance Fields (NeRF)
The Neural Radiance Field (NeRF) [22] establishes a parametric representation of the scene, enabling realistic rendering of novel viewpoints from a given image collection. NeRF is defined by $\mathbf{c},\sigma=F_{\theta}(\mathbf{x},\mathbf{v})$, where $\mathbf{c}$ is output color, $\sigma$ represents opacity, $\mathbf{x}$ is the 3D point position, and $\mathbf{v}$ denotes the viewing direction. The variable $\theta$ represents learnable parameters optimized for each scene.
Considering an input query viewpoint with camera parameters $\mathbf{K}$, $\mathbf{R}$, and $\mathbf{t}$ (representing intrinsic parameters, rotation, and translation), a pixel’s camera viewing ray can be calculated as $\mathbf{r}(t)=\mathbf{o}+t\mathbf{v}$, where $\mathbf{o}$ is the camera center and $\mathbf{v}$ is the ray direction. The pixel color is then computed through volume rendering as $\mathbf{c}(\mathbf{r})=\sum_{i=1}^{N}w_{i}\mathbf{c}_{i}$, with $\mathbf{c}_{i}$ being the NeRF color evaluated from the i-th sampled point $\mathbf{x}_{i}=\mathbf{o}+t_{i}\mathbf{v}$ along the ray. Weight $w_{i}$ is calculated through alpha composition as $w_{i}=\text{exp}\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right)(1-\text{exp}(-\sigma_{i}\delta_{i}))$, where $\sigma_{i}$ defines sampled point’s opacity and $\delta_{i}=t_{i}-t_{i-1}$ defines the distance between the $i$-th sample and its preceding point. The depth of each pixel can also be rendered through volume rendering: $\mathbf{d}(\mathbf{r})=\sum^{n}_{i=1}\hat{w}_{i}t_{i}$, where $\hat{w}_{i}$ is the weight normalized along the ray.
NeRF is trained by minimizing the difference between the rendered pixel value $\mathbf{c}$ and the observed pixel value $\hat{\mathbf{c}}$. We apply per-pixel Huber Loss: $L_{\text{huber}}(\mathbf{c},\hat{\mathbf{c}})$ during the training. Despite its exceptional performance in synthesizing nearby novel views, NeRF falls short in capturing high-quality geometry and rendering extrapolated viewpoints (Figure 6). To overcome these challenges, we propose patch-based supervision and regularization, both in the form of additional loss functions for training NeRF.
@@ 3.2 Distillation of Patch-based Monocular Cues
Learning-based networks for single-image normal and depth prediction can provide robust cues for the geometry of a scene. Taking inspiration from this, prior works [8, 38] exploit monocular depth and normal supervision to determine the geometry of neural fields. However, these works introduce monocular depth supervision on a per-pixel basis, which does not account for that monocular methods tend to only be accurate locally up to a translation and scale. To address this challenge, it is typically necessary to incorporate either per-batch scale estimation or relative depth supervision. We choose to employ a robust scale-and-shift invariant loss function, which provides superior scale invariance.
Given a patch $P=\{p\}$ and the rendered depths $\{{d}_{p}\}$ and normals $\{\mathbf{n}_{p}\}$, our patch-based monocular supervision loss is defined as:
@@
The depth loss $L_{\text{depth}}$ penalizes the discrepancy between the rendered depth $\{{d}_{p}\}$ and the monocular estimation depth $\{\hat{d}_{p}\}$, whereas $L_{\nabla\text{depth}}$ minimizes their gradient discrepancy. Nevertheless, due to the inherent scale and shift ambiguities, directly supervising depth by minimizing the absolute difference will not yield the best results. We instead optimize the difference between depth images with normalized scale and shift:
$L_{\text{depth}}$ $=\sum_{p\subset P}\|\hat{d}_{p}^{{\dagger}}-{d}_{p}\|$ (2)
$L_{\nabla\text{depth}}$ $=\sum_{p\subset P}\|\nabla\hat{d}_{p}^{{\dagger}}-\nabla{d}_{p}\|,$ (3)
where $\hat{d}_{p}^{\dagger}=s\hat{d}_{p}+t$ represents the normalized depth using the optimal scale $s$ and shift $t$ estimated from pixels $\{p\}$ with a least-squares criterion [20], and $\nabla{d}_{p}$ is the gradient value of depth image at pixel $p$.
Our key insight is that having per-patch scale and shift makes the optimization process more robust to the monocular depth local scale drifts.
We also apply a normal loss with the monocular estimation normals $\mathbf{\hat{n}}$ to regularize our rendered density field, following RefNeRF [31]. Like their approach, we define two separate formulations of rendered surface normals. First, we introduce an additional normal rendering MLP head on NeRF. Using this MLP head our NeRF can produce per-ray normal estimation with volume rendering: ${\mathbf{n}}^{\theta}=\sum_{i}w_{i}\mathbf{n}^{\theta}_{i}$. Next, we estimate surface normal as the negative direction of the gradient of the optical density, denoted as $\mathbf{n}^{\nabla}_{i}=-\nabla\sigma_{i}/\|-\nabla\sigma_{i}\|$, with rendered form as ${\mathbf{n}}^{\nabla}=\sum_{i}w_{i}\mathbf{n}^{\nabla}_{i}$. We directly supervise ${\mathbf{n}}^{\theta},{\mathbf{n}}^{\nabla}$ with the monocular surface normal estimate $\hat{\mathbf{n}}$ as:
$\begin{split}L_{\text{normal}}&=\sum_{p\subset P}(1-\cos(\hat{\mathbf{n}}_{p},\mathbf{n}^{\nabla}_{p})+|\hat{\mathbf{n}}_{p}-\mathbf{n}^{\nabla}_{p}|)\\ &+\sum_{p\subset P}(1-\cos(\hat{\mathbf{n}}_{p},\mathbf{n}^{\theta}_{p})+|\hat{\mathbf{n}}_{p}-\mathbf{n}^{\theta}_{p}|)\end{split}$ (4)
Finally, we also apply the gradient loss $L_{\nabla\text{normal}}$ over $\mathbf{n}^{\nabla}$:
$\begin{split}L_{\nabla\text{normal}}&=\sum_{p\subset P}(|\nabla\hat{\mathbf{n}}_{p}-\nabla\mathbf{n}^{\nabla}_{p}|)\end{split}$ (5)

### v2

@@ Abstract
The latest regularized Neural Radiance Field (NeRF) approaches produce poor geometry and view extrapolation for large scale sparse view scenes, such as ETH3D. Density-based approaches tend to be under-constrained, while surface-based approaches tend to miss details. In this paper, we take a density-based approach, sampling patches instead of individual rays to better incorporate monocular depth and normal estimates and patch-based photometric consistency constraints between training views and sampled virtual views. Loosely constraining densities based on estimated depth aligned to sparse points further improves geometric accuracy. While maintaining similar view synthesis quality, our approach significantly improves geometric accuracy on the ETH3D benchmark, e.g. increasing the F1@2cm score by 4x-8x compared to other regularized density-based approaches, with much lower training and inference time than other approaches.
Figure 1: We present MonoPatchNeRF (left) on a large-scale scene facade that contains 68 input views. Our method renders realistic images and accurate normals from the test view and reconstructs the complete mesh compared to baselines [28, 20, 55].
@@ 1 Introduction
Modeling 3D scenes from imagery is useful for mapping, facility assessment, robotics, construction monitoring, and many other applications, which typically require both accurate geometry for measurement and realistic visualization for novel views.
Traditional multi-view stereo (MVS) methods predict accurate geometry given color images, but they have many limitations, such as modeling incomplete surfaces, low ability to render novel views, and reliance on complex algorithms and heuristics that inhibit further improvement. Neural Radiance Field (NeRF) [26] and 3D Gaussian Splatting [14] provide excellent synthesis of novel views, especially when interpolating or near the training views. Recent work aims to improve the geometric accuracy in densely sampled scenes, from object-centric scale [53, 43] to large building scale [55, 20]. However, these approaches do not perform well in large scenes with sparse views (Fig. 1), which is a commonly encountered setting in many applications. In this paper, we aim to create 3D models that provide accurate geometry and view synthesis for large scale scenes with sparse input views.
Current approaches to improve geometric accuracy of NeRF models include guiding with monocular geometry estimates [55], applying appearance priors to virtual views [28], and constraining solutions with SDF-based models [55, 20]. Our early experiments showed SDF-based methods have difficulty capturing details in large, complex scenes, due to sensitivity to initialization, the surface smoothness prior, and limits on volumetric resolution. This leads us to a density-based approach, where the challenge is regularizing an under-constrained solution space.
The key to our approach is to use patch-based ray sampling to provide novel constraints that better incorporate monocular estimates and encourage cross-view consistency. Monocular depth estimators (e.g., [12]) provide excellent estimates of local shape but may not be globally consistent. By sampling multiple rays in a local patch, we can predict depth gradients and enforce consistency up to a scale and translation. Also, rather than a general appearance prior [28], we apply a more specific constraint of occlusion-aware patch-based photometric consistency between sampled virtual views and training views. Density-based models tend to “cheat” by modeling view-dependent effects as densities near the frustrum, which we prevent with loose constraints based on monocular depth maps aligned to sparse structure-from-motion points.
Our experiments show that each of these improvements are critical to improve geometry estimates in large-scale sparse-view scenes. Our method significantly outperforms other NeRF-based approaches on the ETH3D benchmark [36] in geometry estimation while maintaining competitive results in novel view synthesis. We also achieve competitive results on the Tanks and Temples benchmark [15], which has denser views. Our method also has practical advantages of faster training and inference and lower memory requirements, compared to other NeRF-based approaches that aim for geometric accuracy. Our method still falls short of the best classic MVS approaches according to point cloud metrics, but provides a good balance of geometric accuracy and novel view synthesis.
In summary, this paper offers the following main contributions:
  • More effective use of monocular geometry estimates through patch-based ray sampling (Tab. 34)
  • Effective photometric consistency constraints between training and sampled virtual views (Tab. 34)
  • State-of-the-art blend of geometric accuracy and novel view synthesis for complex scenes with sparse views (Tab. 1).
@@ 2 Related Works
Our work aims to create models from sparse views of complex, large-scale scenes that achieve both geometric accuracy, typically pursued by multiview stereo (MVS), and realistic novel view synthesis, as pursued by neural rendering methods.
MVS is a well-studied field, ranging from early works [22] with pure photometric scoring to more recent approaches that incorporate learned features [50, 52, 11, 17]. While scene representations vary, state-of-the-art methods typically predict the depth map of each image based on photometric consistency with a set of source views [34, 16, 23, 17], and use geometric consistency across views to fuse the depths together into a single point cloud [34, 8]. The fused point clouds are evaluated against the ground truth geometry for precision-recall [15], or accuracy-completeness [36], with $F_{1}$ score combining the harmonic mean of the two values. Different from MVS, our method constructs a 3D density representation and extracts depth maps using volumetric rendering of expected depth at source views, followed by a conventional depth-map fusion pipeline [8].
Rather than producing precise 3D points, neural rendering methods, such as Neural Radiance Field (NeRF) [26] and 3D Gaussian Splatting (3DGS) [14], aim to realistically synthesize novel views by optimizing a model to render the training views. Many efforts boost the rendering quality [46, 1, 2], rendering efficiency [27], model size [18, 31], and device requirements [3]. DeLiRa [9] incorporates multiview photometric consistency but notes that this requires overlap between input images, limiting applicability to sparsely viewed scenes.
While NeRF and 3DGS approaches have been shown to reliably estimate geometry and appearance in dense captures [20, 7], multiple papers [54, 13, 30, 28, 48, 42] show failures in cases of outward-facing, wide-baseline, or sparse inputs. PixelNeRF [54] and DietNeRF [13] propose learning based feed-forward solutions that use prior knowledge to better handle sparse inputs. RegNeRF [28] regularizes by applying appearance likelihood and geometric smoothness objectives to patches sampled from virtual views. SPARF [39] jointly optimizes the NeRF models and camera poses with extracted pixel matches on input views, and improves performance given sparse inputs. FreeNeRF [48] improves sparse-view modeling of object-scale scenes by attenuating high-frequency components of the ray positional encoding and penalizing near-field densities. Our strategy is closest to that of RegNeRF [28], as we also sample patches from virtual views to regularize appearance. We enforce a stronger constraint of photometric consistency with training views and also encourage geometric consistency with monocular estimates.
Using image-based priors in regularization has also been shown effective in training neural SDF and NeRF-based models. Roessle et al. [32] and NerfingMVS [45] use a pretrained network to estimate dense depth given sparse SfM points, and supervise the NeRF model with the estimated depth. However, the sparse depth from SfM often contains noise that can be passed down to the dense prediction [4, 29]. NeuralWarp [5] proposes cross-view photometric consistency for training patches as MVS with visibility information from SfM. Our occlusion-aware photometric consistency between virtual and training views takes advantage of the novel view synthesis capability of NeRF to provide more effective constraints for sparse views.
Implicit surface models, such as SDF-based, model the scene in terms of distance to surfaces, rather than point densities. MonoSDF [55] guides signed distance function (SDF) models with monocular estimates of depth and normals, producing accurate surface models from videos of rooms. Neuralangelo [20] improves the ability of SDF-based methods to encode details such as bicycle racks and chair legs using coarse-to-fine optimization and by using numerical gradients to compute higher-order derivatives. However, our experiments show that these methods have limited effectiveness for modeling complex, large-scale scenes from sparse views.
@@ 3 Method
Given a collection of posed images capturing a large scale scene, our goal is to construct a 3D neural radiance field that renders high-quality images and predicts accurate and complete surface geometry. We achieve this by introducing three novel components: 1) patch-based geometry supervision to better leverage the local consistency of monocular cues; 2) patch-based occlusion-aware photometric regularization across virtual and training viewpoints, better guiding NeRF training when input views are sparse; 3) density restriction through monocular geometry estimates and sparse points, inhibiting NeRF from modeling view-dependent effects with densities in unlikely areas. Figure 2 provides a summary of our approach.
3.1 NeRF with Patch-Based Sampling
The Neural Radiance Field (NeRF) [25] establishes a parametric representation of the scene, enabling realistic rendering of novel viewpoints from a given image collection. NeRF is defined by $\mathbf{c},\sigma=F_{\theta}(\mathbf{x},\mathbf{v})$, where $\mathbf{c}$ is output color, $\sigma$ represents opacity, $\mathbf{x}$ is the 3D point position, and $\mathbf{v}$ denotes the viewing direction. The variable $\theta$ represents learnable parameters optimized for each scene. The pixel color and depth can be computed through volume rendering along the corresponding ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{v}$, where $\mathbf{o}$ is the camera center and $\mathbf{v}$ is the ray direction.
NeRF is trained by minimizing the difference between the rendered RGB value $\hat{\mathbf{c}}$ and the observed value $\mathbf{c}$ of random sampled pixels across images. Despite its exceptional performance in synthesizing nearby novel views, NeRF falls short in capturing high-quality geometry and rendering extrapolated viewpoints (Figure 6). To overcome these challenges, we sample local patches instead of discrete rays, and propose our patch-based supervision and regularization. During training, we iterate all training views, sample a batch of local $8\times 8$ patches $P=\{p\}$, and apply the per-pixel Huber loss for each patch: $L_{\text{rgb}}=\sum_{p\subset P}L_{\text{huber}}(({\mathbf{\hat{c}}_{p}},\mathbf{c}_{p}))$.
@@ 3.2 Distillation of Patch-based Monocular Cues
Learning-based networks for single-image normal and depth prediction can provide robust cues for the geometry, especially for areas with photometric ambiguity (reflective and textureless surfaces). Taking inspiration from this, prior works [10, 55] exploit monocular depth and normal supervision for neural fields, and apply a pixel-based scale and shift per batch to align monocular and predicted depth. However, monocular methods tend to be only locally consistent. Image-wide scale-and-shift alignment may still leave large depth errors that reduce the usefulness of monocular depth estimates for supervision. To address this challenge, we compute the scale and shift per local patch to better leverage the capacity of monocular cues and achieve better performance (Compare Mono. and Mono.+Patch in Tables 3 and 4).
For each patch $P=\{p\}$, we compute the transformed monocular depth $\{d_{p}^{\dagger}\}$ using the optimal scale $s$ and shift $t$ from a least-squares criterion [24] between rendered depth$\{\hat{d}_{p}\}$ and monocular depth $\{d_{p}\}$ on $P$. The depth loss $L_{\text{depth}}$ and $L_{\nabla\text{depth}}$ are applied to penalizes the absolute and gradient discrepancy between $\{\hat{d}_{p}\}$ and $\{{d}_{p}^{\dagger}\}$:
$L_{\text{depth}}$ $=\sum_{p\subset P}\|\hat{d}_{p}-{d}_{p}^{{\dagger}}\|$ (1)
$L_{\nabla\text{depth}}$ $=\sum_{p\subset P}\|\nabla\hat{d}_{p}-\nabla{d}_{p}^{{\dagger}}\|,$ (2)
Following RefNeRF [41], we compute density-based normals $\mathbf{n}^{\nabla}_{i}=-\nabla\sigma_{i}/\|-\nabla\sigma_{i}\|$ with the gradient of opacity and MLP-based normals ${\mathbf{n_{i}}}^{\theta}=F_{\theta}(\mathbf{x_{i}},\mathbf{v_{i}})$ with a normals rendering head. After volume rendering, the two normals predictions are supervised with monocular normals $\{{n}_{p}\}$ using angular and $L_{1}$ loss:
$\begin{split}L_{\text{normal}}&=\sum_{p\subset P}(1-\cos({\mathbf{n}}_{p},\mathbf{n}^{\nabla}_{p})+|{\mathbf{n}}_{p}-\mathbf{n}^{\nabla}_{p}|)\\ &+\sum_{p\subset P}(1-\cos({\mathbf{n}}_{p},\mathbf{n}^{\theta}_{p})+|{\mathbf{n}}_{p}-\mathbf{n}^{\theta}_{p}|)\end{split}$ (3)
We also apply the gradient loss $L_{\nabla\text{normal}}$ over $\mathbf{n}^{\nabla}$:
$\begin{split}L_{\nabla\text{normal}}&=\sum_{p\subset P}(|\nabla{\mathbf{n}}_{p}-\nabla\mathbf{n}^{\nabla}_{p}|)\end{split}$ (4)
Finally, our patch-based monocular loss is defined as: